Heat flow through a wooden board 30mm thick at 0.0086w/cm^3 when one of its side 20°C warmer than the other. Determine the thermal conductivity of the wood.
1
Expert's answer
2020-06-30T18:26:28-0400

Solution.

"L=30mm=3cm;"

"\dfrac{Q}{A}=0.0086W/cm^3\sdot3cm=0.0258W/cm^2=258W/m^2;"

"\Delta T=20^oC=20K;"

"\dfrac{\Delta Q}{\Delta t}=" "\dfrac{\Delta T}{L}\sdot K\sdot A";


"\dfrac{Q}{A}=\dfrac{\Delta T}{L}\sdot K\implies K=\dfrac{\ Q}{A}\sdot\dfrac{L}{\Delta T}" ;


"K=258W/m^2\sdot\dfrac{0.03m}{20K}=0.387Js^{-1}m^{-1}K^{-1};"


Answer: "K=0.387JS^{-1}m^{-1}K^{-1}."